OpenStudy (tia_sweetheart17):

solve for x: 8(x-8)+2=-62

OpenStudy (anonymous):

distribute and solve for x 8x-64+2=-62 8x=0 x=0

OpenStudy (amigatour):

8(x-8)+2= -62 Take care of the multiplication and parenthesis first. 8x-64+2 = -62 8x-62 = -62 +62=+62 8x=0 8x/8=0/8 x=0